1. FIND THE DRIFT VELOCITY OF THE FREE ELECTRONS INSIDE THE COPPER METAL WITH AREA 10(-5) METER SQAURE AND CURRENT ACROSS IS 100 A(GIVEN THE DENSITY 8969 KG/ METER CUBE.(can someone explain the equation you use)
1
Expert's answer
2021-10-10T15:56:32-0400

Find the concentration of copper atoms per a cubic meter:


"n=\frac{\rho}{\Mu}N_a=\frac{8969}{0.064}(6.023·10^{23})=8.44·10^{28}\text{ m}^{-1}."

In this equation, NA is Avogadro's number, ρ is density, M is molar mass of copper (64 g/mol).

We know that the current in a conductor can be expressed as


"I=nevA,\\\space\\
v=\frac{I}{neA}=\frac{100}{(8.44·10^{28})(1.602·10^{-19})(10^{-5})}=7.4·10^{-4}\text{ m/s}."

The drift velocity here is around 0.74 mm/s.

In the last equation,

I is the current,

e is the elementary charge,

A is the cross-sectional area of the conductor,

v is the drift velocity of electrons.